by what number must you multiply each side of the equation to produce the equivalent equation x+25?
2/5 x=10
a. 1/5
b. 5/2
c. 4
d. 5

Answer by Nate(3500)

2x/5 = 10 is 2x/5 - 10 = 0
What times -10 gives us 25?
-10x = 25
x = -5/2
Answer: None of the above; I think you typed something wrong.